Help please

Parenteral Dosages Order Atropine salfate 0.15 mg SC sta Supply: Atropine sulfate 0.4 mg per ml. Give mL 2. The drug order reads moephine sulfate gs 1/6 IM q3-4h prn, and the label states morphine sulfate 15 mg per ml. How many mL do you 3. Order. Codeine gr ¼ SC q 4h p.e.a, pain Supply: 20 mL vial Codeine labeled 30 mg per mL. Give: Order: Bicillin 2,400,000 U IM stat Supply: 10 mL. vial of Bicilin containing Give_ 4. 600,000 U per mL 5. Order: Digoxin 600 mcg IV stat Supply: 0.5 mg in 2 mL. Give: mL. 6. Order: Procaine penicillin G 24 million U IM stat Supply: Wycillin (Procaine penicillin G) disposable, single-dose syringe containing 2,400,000 U/2 mL. 7. Order Tigan 200 mg IM stat, then 100 mg q.6h p.t.n. nausea Supply: 2 mL vial Tigan containing 100 mg per mL 8. Order: Heparin 8000 U SC q.8h Supply: 10,000 U per mL. Give_ mL 9. Order: Potassium chloride 15 mEq added to each 1000 mL IV fluid container Supply: Potassium chloride 30 ml. vial containing 2 mEq/mL 10. Order. Demerol 60 mg IM q.4h p.r.n. pain Supply: Demerol 75 mg per 1.5 mL Give mL.
Answer #1

1)0.375 ml of Atropine sulfate,0.4mg=1ml,so ordered dose 0.15 mg=1/0.4x0.15=0.375ml

2)1 mg=1/60 grain(gr),so ordered dose 1/6 gr=10 mg.

now available dose 15mg =1ml.so ordered 10 mg dose=1/15 x10=0.66=0.67mlme 7 Ans 0.5ml-06-Codeine = Order = 2400.000 U So, Ordered dee 너 nl